1. When one junction X of a thermocouple is placed in melting ice and junction Y in steam at 100 degree Celsius the e.m.f is 6.0mV.junction X removed from the melting ice and is placed in a liquid bath at constant temperature, junction Y remaining in steam,e.m.f is now -15mV. What is the temperature of the bath on centigrade scale of this thermocouple?
  2. Find the change in Volume of an aluminum sphere with radius 10cm when the sphere is heated from 0 degree to 💯 degree?
  3. A 32g of water at 60 degree is poured into 60g of cold water at 12degree which is contained in calorimeter of mass 40g and specific heat capacity 0.1J/kg k. Show that the final resulting temperature is 29 degree.
  4. Calculate the root mean square of agas STP if there are 2.69×10power of 23 molecules in 1dm3 and mass of each molecule is 3.35×10-24.
  5. The temperature of a body falls from 30 degree to 20 degree in 5minutes the surrounding air is at 13 degree.find the temperature after further 5minutes.
1
Expert's answer
2021-04-05T11:07:39-0400

1)


"\Delta T_2=\frac{1.5}{6}100=25\degree C\\T_2=100-25=75\degree C"

2)


"\Delta V=\beta V\Delta T\\\Delta V=\frac{4}{3}\pi(10)^3(69\cdot10^{-6})(100)=28.9\ cm^3"

3)


"mc(T_h-T)=mc(T-T_C)+m_cc_c(T-T_C)\\
(32)(4200)(60-T)\\=(60)(4200)(T-12)+(40)(0.1)(T-12)\\T=29\degree C"

4)


"v=\sqrt{\frac{3pV}{Nm}}\\v=\sqrt{\frac{3(101325)(1\cdot10^{-3})}{(2.69\cdot10^{23})(3.35\cdot10^{-24})}}=18.4\frac{m}{s}"

5)


"\frac{\Delta T}{\Delta t}=k(T_h-T_c)\\(T_3-T_2)=(T_2-T_1)\\(T_3-20)=(20-30)\\T_3=10\degree C"